OpenStudy (anonymous):

solve:(x - 3) /x = (x - 3) / (x - 6) im lost with this one can someone show me step by step please

OpenStudy (amistre64):

since the tops are the same; ignore them and equate x with x-6

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

yes, its similar to having 1 over x = 1/x-6

OpenStudy (amistre64):

x = x-6 -x -x -------- 0 = -6 i dont see it haveing a solution tho

OpenStudy (anonymous):

hmmm okay thank you